# Future-Proofing Class Incremental Learning

### v1

@@ 1 Introduction
While humans naturally learn and accumulate knowledge continuously during their whole lifespan, deep neural networks struggle to do the same. Continuously training deep neural networks on new data will inevitably induce a severe decrease of the performance on the previously learned data. This phenomenon, named catastrophic forgetting [29, 38], has received more and more attention during the past decade, and mitigating it is the core problem of incremental learning and continual learning. Among the different settings of incremental learning, Class Incremental Learning (CIL) is among the most challenging ones. In this setting, the model is incrementally trained on a few number of classes at a time, without having access to the previously learned ones. The objective is to train a unified model that performs well on every encountered class.
@@
The rehearsal memory [6, 39] is one of the most commonly used methods to mitigate catastrophic forgetting and is used by most of the approaches for class incremental learning. It consists of a small buffer containing few representative exemplars of each previously learned class that are later used for replay while learning new classes. However, there remain situations where using a rehearsal memory is not possible, either due to privacy concerns regarding the training data or due to limitations in the storage space. This led to the emergence of a new setting named Exemplar-Free Class Incremental Learning (EFCIL) or Class Incremental Learning Without Memory. Methods for EFCIL often freeze the feature extractor after the first incremental step [13, 35], achieving competitive performances while allowing for a faster and less computationally expensive learning of new classes compared to methods that have to backpropagate through the entire deep neural network. However, this may also lead to poor performance when the number of classes available in the initial step to train the feature extractor is too limited.
@@
In this work, we propose to address this limitation by leveraging pre-trained text-to-image diffusion models to better prepare exemplar-free class incremental learning methods relying on a frozen feature extractor for the future. The diffusion models generate synthetic samples of possible future classes that will be jointly used with actual real data during the first incremental step to train a feature extractor that better generalizes to future classes. Experiments using various challenging settings and datasets for EFCIL highlight how our method can be combined with state-of-the-art approaches and results in significant improvement of the average incremental accuracy. Moreover, it appears that using synthetic images of future classes achieves higher performance than using real images from different yet similar classes.
Our contributions can be summarized into two points:
  • 1.
@@
    We propose a framework relying on pre-trained diffusion models to better prepare exemplar-free class incremental learning methods for the future.
  • 2.
@@
    We experimentally show that using synthetic samples of future classes for training the feature extractor achieves higher performance than using real images from different classes.
@@ 2 Related Work
Figure 2: Diagram representing our proposed method Future-Proof Class Incremental Learning (FPCIL) which prepares the feature extractor for the future by jointly training it on an auxiliary dataset in addition to the current dataset during the initial step. The auxiliary dataset contains synthetic samples of future classes generated using a pre-trained text-to-image diffusion model.
2.1 Class Incremental learning (CIL) and exemplar-free class incremental learning (EFCIL)
Most approaches for Class Incremental Learning [3, 28, 60] leverage knowledge distillation [9, 17, 39], a rehearsal memory and a method to mitigate the bias toward the new classes [1, 17, 20, 52] induced by the limited size of the replay memory. Recent works [53, 61] proposed to extend the feature extractor at each incremental step to learn the new classes while preserving past knowledge by freezing previously learned weights.
Exemplar-free class incremental learning is a more challenging setting in which there is no rehearsal memory, making the replay of samples from the previously encountered classes impossible. Similarly to approaches designed for class-incremental learning, numerous methods for exemplar-free class incremental learning also continuously finetune the feature extractor and the classifier [21, 51, 62, 63, 64]. Several authors have highlighted the risk of overfitting on the current task and the importance of having general and transferable feature representation for EFCIL. To address this concern, approaches often rely on self-supervised learning [51, 63] or by creating augmented classes using Mixup [54]. Recently, methods relying on a fixed feature extractor have gained more attention [13, 32]. Those methods either rely on a pre-trained model [45, 49, 50, 57] or by freezing the feature extractor after learning it during the initial step. Petit et al. [35] proposed FeTrIL which achieves state-of-the-art performances by relying on a frozen feature extractor and a linear classifier trained using actual features of new classes and pseudo-features of past classes generated by a geometric translation of new class features. Those methods exhibit numerous advantages including a less computationally expensive training procedure that allows for for a deployment on edge devices. However, they also depend on the initial training strategy used [19, 36] and may achieve poor performance if the quantity of data available for training the feature extractor during the initial step is too limited.
@@ 2.2 Training on synthetic images
While few attempts have been made using GAN-based models [4, 18], the use of synthetic images for training deep learning models has recently gained momentum following the advances in diffusion models [15, 46] and their wide availability. Diffusion models generate synthetic samples using an iterative denoising process conditioned on an input, usually a text prompt. Imagen [42] performs the denoising task in the pixel space while Stable Diffusion [40] and DALL-E2 [37] perform it in the latent space in order to decrease the computational cost of the generative process. In one of the pioneer works, Sariyildiz et al. [43] studied the use of the pre-trained text-to-image diffusion model Stable Diffusion to generate a synthetic replacement for the original training dataset. Their experiments highlighted that while the model trained using the real train dataset reaches a higher accuracy on the real test dataset compared to the one trained on synthetic images, they perform on par for transfer learning tasks. Concurrently, Zhang et al. [56] proposed to expand real datasets using synthetic samples generated with various diffusion models and achieved significant improvement on various image classification datasets. Several authors proposed to use synthetic images generated by pre-trained diffusion models for different problems of computer vision such as zero-shot and few-shot learning [14], long-tailed image classification [44], incremental learning [12, 22], out-of-distribution detection [11], or semantic segmentation [31]. Tian et al. [47] showed that the representation learned using contrastive learning on synthetic images generated by Stable Diffusion outperforms state-of-the-art methods trained on real datasets.
@@ 2.3 Additional data for Incremental Learning
Several authors [2, 25, 26, 27, 55] have considered the use of additional data for class-incremental learning by leveraging datasets of real images belonging to classes different from the ones that the model has to learn. Most of these approaches assume that the additional dataset is unlabeled, restricting the use of those additional data samples to the distillation loss [26, 27, 55]. Jodelet et al. [22] first proposed to use large pre-trained text-to-image diffusion models in order to generate additional synthetic images for the classes previously encountered by the model. Using a textual prompt based on the class name, this method generates labeled samples that can be used for the distillation loss but also for any supervised losses such as the classification loss. Similarly, pre-trained diffusion models were also used for class-incremental semantic segmentation [7] and object detection [23].
@@ 2.4 Preparing for the future
Table 1: Definition of the different auxiliary datasets used during the first incremental step for each setting. “Class overlap” is the percentage of classes (actual number of classes in parentheses) in the auxiliary dataset $\mathcal{S}$ that are also in the target dataset, i.e. that will be encountered during the future incremental steps.
@@
Target dataset Auxiliary dataset $\mathcal{S}$ Nature of data Setting Number of classes $\lvert\mathcal{Y}_{\mathcal{S}}\rvert$ Samples per class $n$ Class overlap with target dataset Note
@@
FPCIL-Partial (k%) Synthetic B50 Inc 50 500 kkof the classes are randomly selected from the target dataset, the remaining are randomly selected from the wrong predictions of GPT-3.5. For k=100, we use the denomination “FPCIL-Oracle”.
@@
FPCIL-GPT Synthetic B0 Inc10 150 500 35.3% (53) Classes the most often predicted by GPT-3.5 based on the classes encountered in the first incremental step.
B0 Inc5 166 500 30.7% (51)
@@
B0 Inc5 102 500 31.4% (32)
FPCIL-GPT R2 Synthetic B0 Inc10 53 500 56.6% (30) Subset of “FPCIL-GPT” containing a more restricted set of classes. More restrictive than R1.
B0 Inc5 58 500 34.5% (20)
ImageNet-Subset (100 classes) ImageNet-Compl. Real B50 Inc 50 $\sim$1300 0% Classes randomly sampled from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
@@
ImageNet-Compl. X2 Real B50 Inc 100 $\sim$1300 0% Classes randomly sampled from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
@@
ImageNet-Compl. Animals Real B0 Inc10 90 $\sim$1300 0% Classes randomly sampled among the “animals” classes from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
@@
FPCIL-Oracle Synthetic B50 Inc 50 1300 100% Future classes perfectly predicted.
B0 Inc10 90 1300 100%
B0 Inc5 95 1300 100%
@@
3 Proposed method
The objective of Class-Incremental Learning (CIL) is to learn a unified model $\theta$ on a gradually increasing set of classes. $\theta$ can be decomposed into $\{\Phi,\mathcal{W}\}$ with $\Phi$ the feature extractor and $\mathcal{W}$ the classifier. The training procedure is divided into $T$ incremental steps, each step consisting of a dataset $\mathcal{D}_{t}$ containing samples from the set of new classes $\mathcal{Y}_{t}$ such that $\forall i,j\in\{1,2,...,T\},\mathcal{Y}_{i}\cap\mathcal{Y}_{j}=\emptyset$ for $i\neq j$. Contrary to standard Class-incremental Learning, in the Exemplar-Free Class Incremental Learning (EFCIL) setting, the model does not have access to any replay memory: during each incremental step $t$, the model $\theta_{t}$ is trained solely on the current dataset $\mathcal{D}_{t}$. After each incremental step, the model is evaluated on the test dataset of all the classes observed so far without having any step descriptor.
3.1 Future-Proof Class Incremental Learning
In order to create a future-proof model, as illustrated in Figure 2, we introduce Future-Proof Class Incremental Learning (FPCIL), a straightforward approach which can be combined with any method for class incremental learning relying on a frozen feature extractor. During the first incremental step, when training the feature extractor, in addition to the initial step dataset $\mathcal{D}_{1}$, an additional dataset $\mathcal{S}$ containing $\lvert\mathcal{Y}_{\mathcal{S}}\rvert$ different classes is used. While any labeled datasets could be used for $\mathcal{S}$, we propose a novel approach which consists in using images from classes that will be encountered by the model during the future incremental steps. In detail, the objective of the initial incremental step is extended from solely learning the initial classes $\mathcal{Y}_{1}$ into learning $\mathcal{Y}_{1}\cup\mathcal{Y}_{\mathcal{S}}$ where $\mathcal{Y}_{\mathcal{S}}=\bigcup_{i=2}^{T}\mathcal{Y}_{i}$ are the future classes. The feature extractor is jointly trained with a classifier on $\mathcal{D}_{1}\cup\mathcal{S}$. At the end of the initial step, the classifier is restricted to the base classes $\mathcal{Y}_{1}$ by dropping the weights corresponding to the future classes from the auxiliary dataset $\mathcal{S}$, and the feature extractor is frozen. Then, the training procedure continues according to the baseline method for the remaining incremental steps.
In the initial step, our proposed method FPCIL trains on both the initial training dataset $\mathcal{D}_{1}$ and an additional dataset $\mathcal{S}$ which contains training data for predicted future classes. As only the data for the current step are available in the Class Incremental Learning setting, we propose to use synthetic data instead of the real ones. We use a pre-trained text-to-image diffusion model to generate synthetic images from the future classes during the initial step. This means that only the labels of the future classes are necessary. We argue that this requirement is practical and is compatible with most real-world applications of class incremental learning. For example, when initially creating a deep learning model, it is common to plan to update the weights later as new data are collected. The nature of the future data is generally known, in particular future classes that may appear, as the data collection process has to be planned and designed in advance due to its high cost in time and resources. More generally, when deploying a learning agent, realistic assumptions on the classes that will be encountered or not can be made as the agent evolves together with the environment; that can be used to prepare the learning agent accordingly.
Based on our experiments detailed in the ablation study, the classes in the auxiliary dataset $\mathcal{S}$ do not need to be perfectly predicted to have a positive impact on the performance. For clarity, we refer to FPCIL as “FPCIL-Oracle” when the list of future classes has been perfectly predicted and “FPCIL-Partial (k%)” otherwise, with $k$ the percentage of future classes correctly predicted.
Predictions of the future classes can be done through expert knowledge or through automated methods. Here, we demonstrate an example of automatically producing predictions for future classes based on the recently popular large language models. Our proposed method relies on the GPT model [5] “GPT-3.5 Turbo Instruct” used through the OpenAI API. The large language model is queried for a completion task using the prompt “The dataset contains the following 100 classes: $c_{1}$, […], $c_{k}$,” where $c_{1}$ to $c_{k}$ are the names of the $k$ classes from the initial step, with $k$ equals to $\lvert\mathcal{Y}_{1}\rvert$. The number “100” is used here to encourage the model to generate a longer list of possible future classes as initial experiments without specifying a number would result in only a few number of guesses. In practice, the output of GPT has to be parsed to extract the predicted future classes. The completion function of GPT often continues the prompt as if it was part of a website or a scientific paper by appending imaginary details such as the number of images per class in the dataset, nonexistent download url for the dataset, or architecture of the model. Due to the random nature of the output from GPT, the prompt is queried 10 times and the predicted classes are ordered according to the number of times they have been predicted. The auxiliary dataset $\mathcal{S}$ is then restricted to the classes which were the most often predicted. We evaluated three versions with an increasing level of restriction: “FPCIL-GPT”, “FPCIL-GPT R1”, and “FPCIL-GPT R2”.

### v2

# Future-Proofing Class-Incremental Learning

@@ 1 Introduction
The version of record of this article, first published in “Machine Vision and Applications”, is available online at Publisher’s website: https://doi.org/10.1007/s00138-024-01635-y
@@
While humans naturally learn and accumulate knowledge continuously during their whole lifespan, deep neural networks struggle to do the same. Continuously training deep neural networks on new data will inevitably induce a severe decrease of the performance on the previously learned data. This phenomenon, named catastrophic forgetting [1, 2], has received more and more attention during the past decade, and mitigating it is the core problem of incremental learning and continual learning. Among the different settings of incremental learning, Class-Incremental Learning (CIL) is among the most challenging ones. In this setting, the model is incrementally trained on a few number of classes at a time, without having access to the previously learned ones. The objective is to train a unified model that performs well on every encountered class.
@@
Rehearsal memory [3, 4, 5] is one of the most commonly used methods to mitigate catastrophic forgetting and is used by most of the approaches for Class-Incremental Learning. It consists of a small buffer containing few representative exemplars of each previously learned class that are later used for replay while learning new classes. However, there remain situations where using a rehearsal memory is not possible, either due to privacy concerns regarding the training data (especially for tasks such as person re-identification) or due to limitations in the storage space. This led to the emergence of a new setting named Exemplar-Free Class-Incremental Learning (EFCIL) or Class-Incremental Learning Without Memory. Methods for EFCIL often freeze the feature extractor after the first incremental step [6, 7], achieving competitive performances while allowing for a faster and less computationally expensive learning of new classes compared to methods that have to backpropagate through the entire deep neural network. However, this may also lead to poor performance when the number of classes available in the initial step to train the feature extractor is too limited.
To address this limitation, several works have proposed to pre-train the feature extractor [5, 8] using additional data from classes similar yet different from those in the target dataset. However, with the advent of foundational diffusion model for image generation, pre-training the feature extractor on a more targeted dataset, specifically classes that are highly likely to be encountered during the future incremental steps, has the potential to lead to better improvements in this setting. While the aspect of using future classes to pre-train the model beforehand seems counter to the concept of Class-Incremental Learning, we argue that it is indeed a realistic setting. In many real world applications, collecting training data for every possible object in an ever changing world to train an omniscient classifier is impossible. It is much more realistic to train a classifier that is more specific to the needs of the user. Even under such a scenario, collecting every possible object that the user needs is an insurmountable task. Therefore, it is much more desirable to attempt to predict specific objects that are likely to be encountered, pre-train the feature extractor, and then perform Incremental Learning while it is deployed to the user. In our case, using a diffusion model that generates training data makes it possible to pre-train especially when objects can be difficult or impossible to obtain.
@@
In this work, we propose to leverage pre-trained text-to-image diffusion models to better prepare Exemplar-Free Class-Incremental Learning methods relying on a frozen feature extractor for the future. The diffusion models generate synthetic samples of possible future classes that will be jointly used with actual real data during the first incremental step to train a feature extractor that better generalizes to future classes. Experiments using various challenging settings and datasets for EFCIL highlight how our method can be combined with state-of-the-art approaches and results in significant improvement of the average incremental accuracy. Moreover, it appears that using synthetic images of future classes achieves higher performance than using real images from different yet similar classes.
Our contributions can be summarized into three points:
@@
  • We propose a framework relying on pre-trained diffusion models to better prepare Exemplar-Free Class-Incremental Learning methods for the future, reaching new state-of-the-art performances.
  • While previous approaches for Incremental Learning used pre-trained diffusion models for generating samples from past classes, our proposed method aims at generating samples from future classes instead.
@@
  • We experimentally show that using synthetic samples of future classes for training the feature extractor achieves higher performance than using real images of classes different from those in the target dataset.
@@ 2 Related Work
2.1 Class-Incremental Learning (CIL) and Exemplar-Free Class-Incremental Learning (EFCIL)
While Class-Incremental Learning may be applied to various tasks and settings such as semantic segmentation [9, 10] or federated learning [11, 12], in this work we only consider the image classification task which is the most extensively studied. Most approaches for Class-Incremental Learning [13, 14] leverage knowledge distillation [3, 15, 16, 17], a rehearsal memory and a method to mitigate the bias toward the new classes [15, 18, 19, 20] induced by the limited size of the replay memory. Recent works [21, 22] proposed to extend the feature extractor at each incremental step to learn the new classes while preserving past knowledge by freezing previously learned weights.
Exemplar-Free Class-Incremental Learning is a more challenging setting in which there is no rehearsal memory, making the replay of samples from the previously encountered classes impossible. Similarly to approaches designed for Class-Incremental Learning, numerous methods for Exemplar-Free Class-Incremental Learning also continuously finetune the feature extractor and the classifier [23, 24, 25, 26, 27]. Several authors have highlighted the risk of overfitting on the current task and the importance of having general and transferable feature representation for EFCIL. To address this concern, approaches often rely on self-supervised learning [23, 26] or by creating augmented classes using Mixup [28]. Recently, methods relying on a fixed feature extractor have gained more attention [6, 29]. Those methods either rely on a pre-trained model [30, 31, 32, 8] or by freezing the feature extractor after learning it during the initial step [6, 7, 33]. Petit et al. [7] proposed FeTrIL which achieves state-of-the-art performances by relying on a frozen feature extractor and a linear classifier trained using actual features of new classes and pseudo-features of past classes generated by a geometric translation of new class features. FeCAM [33] is a recently proposed approach for Exemplar-Free Class-Incremental Learning with a fixed feature extractor that rely on the Mahalanobis distance combined with Tukey’s transformation, covariance shrinkage, and correlation normalization. Those methods exhibit numerous advantages including a less computationally expensive training procedure that allows for for a deployment on edge devices. However, they also depend on the initial training strategy used [34, 35] and may achieve poor performance if the quantity of data available for training the feature extractor during the initial step is too limited.
@@ 2.2 Training on synthetic images
While few attempts have been made using GAN-based models [36, 37], the use of synthetic images for training deep learning models has recently gained momentum following the advances in diffusion models [38, 39] and their wide availability. Diffusion models generate synthetic samples using an iterative denoising process conditioned on an input, usually a text prompt. Imagen [40] performs the denoising task in the pixel space while Stable Diffusion [41] and DALL-E2 [42] perform it in the latent space in order to decrease the computational cost of the generative process. In one of the pioneer works, Sarıyıldız et al. [43] studied the use of the pre-trained text-to-image diffusion model Stable Diffusion to generate a synthetic replacement for the original training dataset. Their experiments highlighted that while the model trained using the real train dataset reaches a higher accuracy on the real test dataset compared to the one trained on synthetic images, they perform on par for transfer learning tasks. Concurrently, Zhang et al. [44] proposed to expand real datasets using synthetic samples generated with various diffusion models and achieved significant improvement on various image classification datasets. Several authors proposed to use synthetic images generated by pre-trained diffusion models for different problems of computer vision such as zero-shot and few-shot learning [45], long-tailed image classification [46], incremental learning [47, 48], out-of-distribution detection [49], or semantic segmentation [50]. Tian et al. [51] showed that the representation learned using contrastive learning on synthetic images generated by Stable Diffusion outperforms state-of-the-art methods trained on real datasets.
@@ 2.3 Additional data for Incremental Learning
Several authors [52, 53, 54, 55, 56] have considered the use of additional data for Class-Incremental Learning by leveraging datasets of real images belonging to classes different from the ones that the model has to learn. Most of these approaches assume that the additional dataset is unlabeled, restricting the use of those additional data samples to the distillation loss [52, 53, 56]. Jodelet et al. [47] first proposed to use large pre-trained text-to-image diffusion models for Incremental Learning in order to generate additional synthetic images for the classes previously encountered by the model. Using a textual prompt based on the class name, this method generates labeled samples that can be used for the distillation loss but also for any supervised losses such as the classification loss. More recently, pre-trained diffusion models were also used for Class-Incremental semantic segmentation [57] and object detection [58]. However, there remains a significant distribution gap between the real samples and the synthetic ones [47] whose impact on the classifier makes the use of pre-trained diffusion model for replaying samples from past classes especially difficult for Exemplar-Free Class-Incremental Learning.
@@ 2.4 Preparing for the future
3 Proposed method
The objective of Class-Incremental Learning (CIL) is to learn a unified model $\theta$ on a gradually increasing set of classes. $\theta$ can be decomposed into $\{\Phi,\mathcal{W}\}$ with $\Phi$ the feature extractor and $\mathcal{W}$ the classifier. The training procedure is divided into $T$ incremental steps, each step consisting of a dataset $\mathcal{D}_{t}$ containing samples from the set of new classes $\mathcal{Y}_{t}$ such that $\forall i,j\in\{1,2,...,T\},\mathcal{Y}_{i}\cap\mathcal{Y}_{j}=\emptyset$ for $i\neq j$. Contrary to standard Class-Incremental Learning, in the Exemplar-Free Class-Incremental Learning (EFCIL) setting, the model does not have access to any replay memory: during each incremental step $t$, the model $\theta_{t}$ is trained solely on the current dataset $\mathcal{D}_{t}$. After each incremental step, the model is evaluated on the test dataset of all the classes observed so far without having any step descriptor.
3.1 Future-Proofing Class-Incremental Learning
Figure 1: Diagram representing our proposed method Future-Proof Class Incremental Learning (FPCIL) which prepares the feature extractor for the future by jointly training it on an auxiliary dataset in addition to the current dataset during the initial step. The auxiliary dataset contains synthetic samples of future classes generated using a pre-trained text-to-image diffusion model. After the first incremental step, the feature extractor $\phi$ is frozen and the weights of the classifier $\tilde{w}$ corresponding to the future classes are removed. Afterward, in the following incremental step, only the classifier $\mathcal{W}$ is trained using dedicated method (e.g. FeTrIL, FeCAM).
In order to create a future-proof model, as illustrated in Figure 1, we introduce Future-Proofing Class Incremental Learning (FPCIL), a straightforward approach which can be combined with any method for Class-Incremental Learning relying on a frozen feature extractor. During the first incremental step, when training the feature extractor, in addition to the initial step dataset $\mathcal{D}_{1}$, an additional dataset $\mathcal{S}$ containing $\lvert\mathcal{Y}_{\mathcal{S}}\rvert$ different classes is used. While any labeled datasets could be used for $\mathcal{S}$, we propose a novel approach which consists in using images from classes that will be encountered by the model during the future incremental steps. In detail, the objective of the initial incremental step is extended from solely learning the initial classes $\mathcal{Y}_{1}$ into learning $\mathcal{Y}_{1}\cup\mathcal{Y}_{\mathcal{S}}$ where $\mathcal{Y}_{\mathcal{S}}=\bigcup_{i=2}^{T}\mathcal{Y}_{i}$ are the future classes. The feature extractor is jointly trained with a classifier on $\mathcal{D}_{1}\cup\mathcal{S}$. At the end of the initial step, the classifier is restricted to the base classes $\mathcal{Y}_{1}$ by dropping the weights corresponding to the future classes from the auxiliary dataset $\mathcal{S}$, and the feature extractor is frozen. Then, the training procedure continues according to the baseline method for the remaining incremental steps. Our proposed method is detailed in Algorithm 1.
Input: Data-flow $\{\mathcal{D}_{i}\}_{i=1}^{T}$ and pre-trained generative model G.
Ouput: Models $\{\theta_{i}\}_{i=1}^{T}$ composed of $\{\Phi_{i},\mathcal{W}_{i}\}$.
for $t\in\{1,2,...,T\}$ do
       if $t=1$ then
             // Predict future classes based on initial classes using oracle or GPT
             $\mathcal{Y}_{\mathcal{S}}\leftarrow$ PredictFuture($\mathcal{Y}_{1}$)
             // Generate synthetic samples of future classes
             $\mathcal{S}\leftarrow$ GenerateSynthetic(G, $\mathcal{Y}_{\mathcal{S}}$)
            // Train the model using real data of the first step and synthetic data of future steps
             Train $\theta_{1}$ using $\mathcal{D}_{1}\cup\mathcal{S}$
             Freeze $\Phi_{1}$ and remove $\mathcal{Y}_{\mathcal{S}}$ weights from $\mathcal{W}_{1}$
      else
             // Incrementally update the classifier using dedicated method (e.g. FeTrIL, FeCAM)
             Train $\mathcal{W}_{t}$ using $\mathcal{D}_{t}$
Algorithm 1 Future-Proof Class-Incremental Learning (FPCIL)
In the initial step, our proposed method FPCIL trains on both the initial training dataset $\mathcal{D}_{1}$ and an additional dataset $\mathcal{S}$ which contains training data for predicted future classes. As only the data for the current step are available in the Class-Incremental Learning setting, we propose to use synthetic data instead of the real ones. We use a pre-trained text-to-image diffusion model to generate synthetic images from the future classes during the initial step. This means that only the labels of the future classes are necessary. Compared to recent approaches for Incremental Learning relying on a pre-trained diffusion model [47, 57, 58] for generating synthetic samples of past classes, our proposed method aims at generating synthetic samples of future classes instead. By primarily using the synthetic data to train the feature extractor, our proposed method circumvents the impact of the distribution gap between real and synthetic samples generated by the diffusion model, making it applicable to Exemplar-Free Class-Incremental Learning.
While the requirement of having access to the label of future classes may seem counter to the concept of Class-Incremental Learning, we argue that it is practical and compatible with most real-world applications of Class-Incremental Learning. In general, during the developing a deep learning model meant for continuously accumulating new knowledge while achieving a given task in an ever-evolving environment, both the task and the environment are known. Therefore, it is reasonable to assume that it is possible, based on this initial knowledge of the environment and task, to predict their possible future evolution. However, acquiring beforehand real data corresponding to these future states of the environment and task may be difficult, expensive, or even impossible in some cases which lead us to use synthetic samples instead. As a concrete example, we propose to consider the case of an agent that would be deployed on an embedded device for assisting its user inside his home. The agent would have to be able to learn new concepts based on data directly collected from its environment depending on the particular need of its user. It would have to rely on incremental learning as frequent retraining from scratch would not be possible on either a cloud server as uploading user’s data would be difficult due to privacy reasons, or in local due to the limited computational and memory capacity of the embedded device. Therefore, using our proposed method it is possible to train an initial model precisely tailored for each user (or group of users) before deployment, by relying on available knowledge about the user and its environment (interests, location, particular requests, etc.) to preemptively learn specific concepts that may appear in the future in addition to generic concepts. Since building a specific dataset of real samples for all possible needs of all possible users would not be tractable due to the significant cost in time and resources of the data collection process, the use of synthetic data appears as a faster and cheaper alternative. As a second example, we consider the case of an agent deployed on an embedded device for performing a task in a hard to access and potentially hostile environment where the agent has to be autonomous due to unreliable remote control possibility (deep sea, space, etc). The nature of the environment intrinsically limits the capacity to collect large training datasets in advance but its properties may be known. Using our proposed method, the agent could be trained before its deployment using synthetic data from generative diffusion models, or more realistically in this scenario using simulation-based tools. The agent would then be able to incrementally learn and adjust itself to the environment using real data collected while performing the task.
It should also be noted that based on our experiments detailed in the ablation study, the classes in the auxiliary dataset $\mathcal{S}$ do not need to be perfectly predicted to have a positive impact on the performance. For clarity, we refer to FPCIL as “FPCIL-Oracle” when the list of future classes has been perfectly predicted and “FPCIL-Partial (k%)” otherwise, with $k$ the percentage of future classes correctly predicted.
Predictions of the future classes can be done through expert knowledge or through automated methods. Here, we introduce a simple approach for automatically predicting the future classes based on recently popular large language models. Our proposed method relies on the GPT model [62] “GPT-3.5 Turbo Instruct” used through the OpenAI API. The large language model is queried for a completion task using the prompt “The dataset contains the following 100 classes: $c_{1}$, […], $c_{k}$,” where $c_{1}$ to $c_{k}$ are the names of the $k$ classes from the initial step, with $k$ equals to $\lvert\mathcal{Y}_{1}\rvert$. The number “100” is used here to encourage the model to generate a longer list of possible future classes as initial experiments without specifying a number would result in only a few number of guesses. In practice, the output of GPT has to be parsed to extract the predicted future classes. The completion function of GPT often continues the prompt as if it was part of a website or a scientific paper by appending imaginary details such as the number of images per class in the dataset, nonexistent download url for the dataset, or architecture of the model. Due to the random nature of the output from GPT, the prompt is queried 10 times and the predicted classes are ordered according to the number of times they have been predicted. The auxiliary dataset $\mathcal{S}$ is then restricted to the classes which were the most often predicted. We evaluated three versions with an increasing level of restriction: “FPCIL-GPT”, “FPCIL-GPT R1”, and “FPCIL-GPT R2”. Those approaches based on GPT also serve to imitate a more realistic application of Future-Proof Class-Incremental Learning by predicting various number of future classes with various accuracy.
3.2 Generating future data
Leveraging recent advances in diffusion models and the wide availability of generative text-to-image models pre-trained on natural images, we propose to use them for generating the synthetic images dataset of future classes following the procedure initially defined by Sarıyıldız et al. [43].
The auxiliary dataset $\mathcal{S}$ is generated using a pre-trained diffusion model $G$ conditioned solely on a textual prompt. As proposed in [43], the default prompt used is $c,d_{c}$ with $c$ the name of the class and $d_{c}$ a simple description of the class. Adding the description of the class into the prompt is intended to ensure that the generated images belong to the targeted class and not to any homonyms: for example “crane” in the ImageNet dataset, which may refer to either the bird or the machine used to move heavy objects. The definition $d_{c}$ also tends to restrain the domain of the generated images even though the domain could also be enforced by adding it to the prompt. To automatically create the prompts, we rely on WordNet [63] and use the lemmas of the synset as $c$ and the definition of the synset as $d_{c}$. For each class, $n$ synthetic images are generated with $n$ equals to the number of images per class in the target dataset. Compared to traditional pre-training and concurrent method [55], our proposed method does not require the costly and fastidious process of collecting, curating, and annotating a dataset of real images. Examples of synthetic images generated for CIFAR100 are displayed in Figure 2.
While this generation procedure allows our proposed method to achieves significant improvement compared to baselines, there remains an important distribution gap between the real and synthetic data [47]. Reducing this distribution gap and its impacts on the training of the feature extractor would further improve the performance of our proposed method. As possible future research direction, the diffusion model could be finetuned using the data from the first incremental step, a more sophisticated prompting mechanism could be implemented, or a selection process could be used to discard the synthetic samples too far away from the distribution of the target dataset.
@@ 3.2 Generating future data
Table 1: Definition of the different auxiliary datasets used during the first incremental step for each setting. “Class overlap” is the percentage of classes (actual number of classes in parentheses) from the future incremental steps of the target dataset that are in the auxiliary dataset $\mathcal{S}$, i.e. the number of correctly predicted future classes.
@@
Target dataset Auxiliary dataset $\mathcal{S}$ Nature of data Setting Number of classes $\lvert\mathcal{Y}_{\mathcal{S}}\rvert$ Samples per class $n$ Class overlap with future steps of target dataset Note
@@
FPCIL-Partial (k%) Synthetic B50 Inc 50 500 kkof the classes have been randomly selected from the target dataset, the remaining have been randomly selected from the wrong predictions of GPT-3.5. For k=100, we use the denomination “FPCIL-Oracle”.
@@
FPCIL-GPT Synthetic B0 Inc10 150 500 58.9% (53) Classes the most often predicted by GPT-3.5 based on the classes encountered in the first incremental step.
B0 Inc5 166 500 53.7% (51)
@@
B0 Inc5 102 500 33.7% (32)
FPCIL-GPT R2 Synthetic B0 Inc10 53 500 33.3% (30) Subset of “FPCIL-GPT” containing a more restricted set of classes. More restrictive than R1.
B0 Inc5 58 500 21.0% (20)
ImageNet-Subset (100 classes) ImageNet-Compl. Real B50 Inc 50 $\sim$1300 0% Classes have been randomly sampled from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
@@
ImageNet-Compl. X2 Real B50 Inc 100 $\sim$1300 0% Classes have been randomly sampled from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
@@
ImageNet-Compl. Animals Real B0 Inc10 90 $\sim$1300 0% Classes have been randomly sampled among the “animals” classes from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
@@
FPCIL-Oracle Synthetic B50 Inc 50 1300 100% (50) Future classes perfectly predicted.
B0 Inc10 90 1300 100% (90)
B0 Inc5 95 1300 100% (95)
@@
4 Experiments
4.1 Experimental setup
4.1.1 Datasets
The experiments are conducted on two large scale datasets: CIFAR100 [65] and ImageNet-Subset. CIFAR100 is a dataset containing 60,000 32$\times$32 RGB images evenly divided among 100 classes with 500 training images and 100 testing images per class. ImageNet-Subset is subset of the high-resolution images dataset ImageNet (ILSVRC 2012) [66]. This subset contains 100 randomly selected classes with about 1,300 training images and 50 testing images per class. Following Rebuffi et al. [3], the classes are ordered using numpy with the random seed 1993.
4.1.2 Model training
Following common practice for Exemplar-Free Class-Incremental Learning [7, 26, 27], ResNet-18 is used for both CIFAR100 and ImageNet-Subset. The feature extractor is trained solely during the first incremental step and is frozen for the rest of the training procedure. The backbone is trained with SDG using the standard Softmax Cross-Entropy loss and AutoAugment [67] for 160 epochs with a batch size of 128 and a weight decay of $5e^{-4}$. The initial learning rate is set to 0.1 and decrease to zero using a cosine annealing schedule. For the classifier, we use a fully-connected linear layer trained following FeTrIL [7]. The classifier is trained using SGD for 50 epochs at each incremental step, with an initial learning rate of 0.1 and cosine annealing schedule. In the ablation study, we also report results using a Nearest Class Mean (NCM) Classifier and FeCAM [33]. If not stated otherwise, the reported results for the different baselines have been reproduced by our own experiments. Experiments are conducted using Pytorch [68] and PyCIL [69].